A coin placed 29.0 cm from the center of a rotating, horizontal turntable slips when its speed is 50.2 cm/s.

(a) What force causes the centripetal acceleration when the coin is stationary relative to the turntable?
static frictioninertia    kinetic frictioncentrifugal force

(b) What is the coefficient of static friction between coin and turntable?
Answer #1

a) static friction

centripetal force = (m*v^2)/r Static Friction Force= uS*Fn (the normal force)


b) F(friction) = centripetal force = m*(.502^2)/.29 = uS*m*g

uS:coefficient of static friction m:mass of coin

uS= (.502^2)/(.29(9.8))

=0.0866
Answer #2

a) static friction

centripetal force = (m*v^2)/r Static Friction Force= uS*Fn (the normal force)


b) F(friction) = centripetal force = m*(.502^2)/.299 = uS*m*g

uS:coefficient of static friction m:mass of coin

uS= ((.502)^2)/(.299(9.8))=0.086
Answer #3

a) Force is the sideway frictional force of the coin and table .

b) Friction = mV